Psychometric scores for anxiety, depression, negative self, somatization, and hostility were combined into a single Global Severity Index (GSI). Global Severity Index was compared for 65 adolescents who were seen in the outpatient clinic for 7 different physical complaints (short stature, obesity, lack of weight gain, breast problems, genital problems, menstrual problems, and hirsutism).

(a) Identify the explanatory variable from the above description

psychometric scores

physical complaints     

short stature

Global Severity Index

(b) Identify the response variable from the above description

psychometric scores

physical complaints     

short stature

Global Severity Index

Report the degrees of freedom for the relevant F statistic in the numerator and the denominator. (Round your answers to the nearest whole number.)
Answer #1

a) Explanatory Variable:

physical complaints     

b) Response variable:

Global Severity Index

c) Numerator df = 7 - 1 = 6

Denominator df = 65 - 7 = 58